### v1

Abstract
[…]
[name=Andrea Esposito, color=red]AE abstract The integration of Artificial Intelligence (AI) in modern society is heavily shifting the way that individuals carry out their tasks and activities. Employing AI-based systems raises challenges that designers and developers must address to ensure that humans remain in control of the interaction process, particularly in high-risk domains. This article presents a novel End-User Development (EUD) approach for black-box AI models through a redesigned user interface in the Rhino-Cyt platform, a medical AI-based decision-support system for medical professionals (more precisely, rhinocytologists) to carry out cell classification. The proposed interface empowers users to intervene in AI decision-making process by editing explanations and reconfiguring the model, influencing its future predictions. This work contributes to Human-Centered Artificial Intelligence (HCAI) and EUD by discussing how explanation-driven interventions allow a blend of explainability, user intervention, and model reconfiguration, fostering a symbiosis between humans and user-tailored AI systems.
Keywords:
Customization Black-box AI Model Reconfiguration Explainable AI (XAI) Human-AI symbiosis
abstract The integration of Artificial Intelligence (AI) in modern society is heavily shifting the way that individuals carry out their tasks and activities. Employing AI-based systems raises challenges that designers and developers must address to ensure that humans remain in control of the interaction process, particularly in high-risk domains. This article presents a novel End-User Development (EUD) approach for black-box AI models through a redesigned user interface in the Rhino-Cyt platform, a medical AI-based decision-support system for medical professionals (more precisely, rhinocytologists) to carry out cell classification. The proposed interface empowers users to intervene in AI decision-making process by editing explanations and reconfiguring the model, influencing its future predictions. This work contributes to Human-Centered Artificial Intelligence (HCAI) and EUD by discussing how explanation-driven interventions allow a blend of explainability, user intervention, and model reconfiguration, fostering a symbiosis between humans and user-tailored AI systems.
EUD
End-User Development
AI
Artificial Intelligence
HCAI
Human-Centered Artificial Intelligence
XAI
eXplainable Artificial Intelligence
SAI
Symbiotic Artificial Intelligence
UI
User Interface
1 Introduction
AI has become an integral component of decision-support systems in numerous domains, including healthcare, finance, and law [15]. While AI models could increase efficiency and decision-making capabilities, their reliance on complex, often opaque, algorithms presents a significant barrier to adoption [15]. In high-stakes applications such as medical diagnostics, end-users—typically domain experts rather than AI specialists—require mechanisms to refine and adjust AI behavior to better align with their expertise and contextual knowledge [8]. Traditionally, AI systems have followed a one-size-fits-all approach, where models are trained on large datasets but offer limited opportunities for users to modify their behavior post-deployment [7]. This lack of adaptability can lead to misaligned recommendations, loss of trust, and decreased usability. EUD for AI seeks to address this challenge by enabling non-technical users to customize AI behavior according to their specific needs and preferences [12]. However, existing EUD approaches primarily focus on low-code or no-code AI development environments, offering component-based or rule-based interactions [18]. Few solutions have explored explanation-driven interventions, where users influence AI behavior by modifying its justifications rather than its internal mechanics. The integration of EUD for AI has made substantial progress in fields such as the Internet of Things (IoT), education, and business analytics [11, 18]. However, AI-based decision-support systems, especially those powered by black-box models, remain challenging for users to directly intervene on. Current approaches to AI customization for end-users are (i) Rule-Based Customization[11], (ii) Low-Code / No-Code AI[18], and (iii) Human-AI Collaboration Interfaces[10, 20]. The first one involves the definition from the user of if-this-then-that conditions to influence AI outputs. Although effective in structured environments, this approach lacks flexibility in complex decision-making scenarios. On the other hand, the platforms that adopt Low-Code / No-Code AI allow users to build and deploy AI models without programming (e.g., AutoML tools), but they do not enable real-time intervention on model behavior post-deployment. Lastly, Human-AI Collaboration Interfaces enable users to validate and/or override the system’s predictions without affecting its future behavior. These approaches do not fully address the need for an interactive and iterative refinement of AI behavior based on human expertise, which can be highly useful in critical fields. For example, medical professionals might encounter AI misclassifications that can be reviewed and corrected through targeted feedback. Although the latter is usually employed to merely accept or reject the output without repercussions on the model, using professionals’ knowledge and expertise to refine AI reasoning can be a valuable resource to improve the system’s performance, increase accuracy for future predictions and build a stronger symbiotic relationship between the two parties [14, 10]. Creating AI-based systems that embody these characteristics can foster collaboration, which is central in high-stakes domains. Thus, end-users must be provided with clear, appropriate, and effective interaction mechanisms that enable bidirectional communication, establishing a symbiotic relationship between humans and AI. Symbiotic Artificial Intelligence (SAI) is a specialization of Human-Centered AI [24] and aims at supporting humans instead of replacing them. This implies creating solutions that reflect humans’ needs and preferences by integrating intervention paradigms, transparency, and fairness by design focusing on augmentation rather than automation [16, 10]. This research presents a novel intervention-based User Interface (UI) within the Rhino-Cyt platform, designed to empower rhinocytologists to modify AI-generated classifications and explanations. Rhino-Cyt is an AI-assisted EUD environment for the classification of nasal cytology samples that supports its end-user developers, i.e., medical professionals, in diagnosing inflammatory and allergic conditions [2]. Rhino-Cyt aims at embodying the EUDability construct [5] that introduces an innovative explanation-driven EUD approach, allowing end-users to adjust AI classifications, edit AI-generated explanations, and indirectly refine and tailor the AI model through the mechanism of interventions [22]. This approach goes beyond rule-based or component-based customization, offering a human-in-the-loop model refinement paradigm. Thus, Rhino-Cyt involves rhinocytologists as its end-user developers, supporting them in reaching the goal of diagnosing. The rest of the article is structured as follows. 2 discusses prior research in EUD for AI, explainability, and human-AI collaboration, positioning Rhino-Cyt within this landscape. 3 presents the design of the intervention-based UI, detailing its interaction flow and impact on AI adaptation. 4 positions Rhino-Cyt within existing EUD for AI taxonomies and compares it with other customization paradigms. 5 concludes this article by summarizing its key contributions and outlining the next steps.
2 Background and Related Work
This section reviews prior research on EUD for AI, explainability as a mechanism for EUD, and human-AI collaboration in decision-support systems. In the end, we highlight the explanation-driven intervention paradigm for AI-assisted medical diagnostics as the main contribution of this work.
2.1 End-User Development for AI
EUD encompasses a range of methods, techniques, tools, and socio-technical environments that empower non-professionals to engage in activities usually reserved for professionals in ICT-related areas, including the ability to create, modify, extend, and test digital artifacts without requiring specialized knowledge in conventional software engineering practices [4]. In the context of AI, EUD plays a crucial role in enabling non-technical users to adjust AI systems without requiring programming expertise. A systematic literature review by Esposito et al. categorized existing EUD for AI approaches into five key paradigms [11]:
  1. 1.
    Component-Based: Users assemble predefined AI components through visual programming interfaces.
  2. 2.
    Rule-Based: Users are allowed to modify AI behavior through “if-then” rules.
  3. 3.
    Wizard-Based: Step-by-step guidance simplifies AI customization, presenting the task as a sequence of operations that guide users throughout the overall activity.
  4. 4.
    Template-Based: Users adjust pre-built AI models by modifying parameters.
  5. 5.
    Workflow and Data Diagrams: Users define AI processes using flow-based representations.
Most EUD for AI systems fall within component-based and rule-based paradigms, where users interact with structured representations of AI logic [11]. While these approaches are effective for tasks such as building AI models from scratch or configuring predefined automation rules, they offer limited support for modifying existing black-box AI models. The Rhino-Cyt intervention interface extends EUD for AI by introducing a new paradigm: explanation-driven interventions. Instead of requiring users to manipulate AI model components or logic directly, this approach allows them to edit AI-generated explanations, indirectly refining the model’s behavior over time. This method aligns with the goal of making AI more adaptable to domain-specific knowledge while minimizing technical barriers.
2.2 Explainability as a Mechanism for End-User Development
[…]
XAI seeks to make AI model decisions more interpretable and transparent, particularly for domain experts who rely on AI assistance in high-stakes decision-making scenarios [11, 21]. Traditionally, eXplainable Artificial Intelligence (XAI) was mostly used for post-hoc justification, trust calibration, and bias detection. Through XAI, users can understand AI decisions, deciding whether to rely on its predictions or potentially recognize (and mitigate) biases by assessing its reasoning [15]. However, explainability has rarely been explored as an active mechanism for EUD. In most AI-assisted decision systems, explanations are static: users can view and interpret them, but they cannot modify them to influence future AI behavior. Rhino-Cyt introduces a novel editable explanation mechanism where users can: modify explanations associated with AI classifications, provide domain-specific refinements to ensure AI-generated explanations align with expert knowledge, and influence future AI behavior indirectly, fostering an interactive trust-building process.
2.3 Human-AI Collaboration in Decision Support Systems
AI-based decision-support systems, especially in medicine, often follow a human-on-the-loop paradigm, where users oversee the decision-making process by interacting with AI outputs to merely validate its decisions [23, 10, 13]. This translates into a validation-based collaboration, where experts review AI predictions but have no direct manipulation mechanism for modifying the AI’s reasoning process [10]. Rhino-Cyt aims at filling this gap following the model of human–AI interaction proposed by Desolda et al. [9], moving beyond this by enabling direct, explanation-driven interventions, ensuring that AI’s decision-making processes evolve alongside domain experts, fostering a symbiosis among humans and AI [10]. Establishing proper collaboration mechanisms between humans and AI is crucial to guarantee that professionals are aware of the processes that lead to outputs; in fact, the outcome of the interaction must be a fair, non-discriminatory, and unbiased decision that minimizes potentially negative repercussions and holds accountable both parties. At the same time, humans remain the ones responsible for perpetrating the decision in their application domain [25, 6].
3 Design of the Intervention-Based User Interface
(a) The intervention-based UI
(b) Key areas of the UI
Figure 1: The user interface for allowing black-box AI model tailoring implemented in Rhino-Cyt
The Rhino-Cyt platform introduces an innovative intervention-based UI that allows rhinocytologists to refine AI-assisted nasal cytology classifications, presented in 1. Unlike conventional AI customization methods, which rely on rule-based or low-code paradigms, Rhino-Cyt allows users to intervene in AI decisions and modify AI-generated explanations, indirectly influencing the model’s future behavior [22]. This section details the design principles, UI components, and interaction workflow, demonstrating how the system aligns with EUD for AI by offering an explanation-driven intervention mechanism.
3.1 Design Rationale and Principles
The design of the Rhino-Cyt intervention UI follows three core principles derived from EUD for AI research [11], which emphasize the human intervention that leads to the AI customization while minimizing humans’ cognitive load. The referred principles are:
  1. 1.
    Intervention-Based Interaction. The system enables users to adjust AI-generated classifications and explanations in a structured manner. Any modifications made by users are systematically logged and tracked, allowing for a comprehensive assessment of their impact on the AI model’s behavior over time.
  2. 2.
    Explainability-Driven Customization. The system supports explainability-driven customization by allowing users to refine AI-generated justifications without requiring them to modify raw model parameters or write explicit rules. Instead, the interface leverages editable explanations as a means of customization, enabling domain experts to adjust and improve AI-generated reasoning based on their medical expertise.
  3. 3.
    Minimal Cognitive Load for Domain Experts. The system is designed to minimize the cognitive load for domain experts by providing guided interventions that simplify the interaction process and reduce the need for technical expertise. AI adaptation occurs implicitly through user feedback, allowing the model to refine its reasoning without requiring manual retraining.
3.2 User Interface Components
The elements of Rhino-Cyt’s UI that allow users to intervene on AI decisions consist of three key components:
  1. 1.
    The Classification Details and Interventions Panel presents AI-generated classifications for nasal cytology samples, allowing users to review system decisions. If necessary, users can override AI’s decisions by selecting an alternative category, ensuring that domain expert judgment remains central to the decision-making process. Any corrections made by users are logged and incorporated into the system, contributing to the continuous adaptation and improvement of the AI’s future performance.
  2. 2.
    The Editable Explanation Area displays the AI-generated rationale behind its classifications, providing transparency into the decision-making process. Users can modify these explanations to guarantee they align more closely with medical reasoning and domain expertise. Any edits made by users directly influence the AI’s reasoning model, refining its approach and shaping future justifications.
  3. 3.
    The Impact Visualization Dashboard provides a summary of user interventions, detailing key aspects such as the number of classification overrides, the frequency and nature of explanation edits, and the ways in which AI predictions evolve over time based on these interventions. This feature plays a crucial role in trust calibration by offering users insights into how the AI adapts and improves, helping users to understand how AI behavior is adapting.
By combining those, Rhino-Cyt enables domain experts to refine the AI decision-making process by directly manipulating its output, without requiring technical expertise.
3.3 Interaction Workflow: How Users Intervene in AI Decisions
The Rhino-Cyt intervention workflow is structured as a three-step process, ensuring smooth user interaction with the AI system.
  1. Step 1:
    Reviewing AI-Generated Classification and Explanation. Upon analyzing a nasal cytology sample, the AI presents a predicted classification along with a textual explanation justifying the classification.
  2. Step 2:
    User Intervention via Adjustment or Explanation Editing. The users have two options for intervention: they can either override the AI classification by selecting an alternative label or modify the AI-generated explanation to reflect expert reasoning more accurately, or both. Any changes are recorded, with an optional comment field for contextualizing edits.
  3. Step 3:
    Model Adaptation and Visualization of Impact. The system logs interventions and updates the AI’s explanation model allowing users to track how their interventions can shape future AI predictions.
This workflow enables AI adaptation to be progressive, ensuring the model evolves alongside experts’ knowledge, thus creating a human-in-the-loop customization mechanism for AI-supported decision-making.
3.4 Underlying AI Model Adaptation
Rhino-Cyt employs a hybrid adaptation mechanism. It balances manual user intervention with an automated model refinement process based on users’ feedback, thus providing an example of both an adaptable and adaptive system [12]. The system incremental adaptation is powered by three key requirements: (i) interaction logging, (ii) a quick AI training loop, and (iii) a trust-calibration mechanism. The platform continuously logs user interactions, capturing every action performed by users. This includes explicit feedback on the AI model’s predictions, where users can either accept the suggested classification or reject it and make modifications. This process is facilitated through the “3. Intervention” panel, illustrated in 1(b). The collected feedback is then leveraged to initiate the retraining of the AI model. To enhance the adaptivity and adaptability of the AI system, several techniques can be employed, ranging from online learning [1, 26] to reinforcement learning from human feedback [17], ensuring timely updates to the model. However, effective adaptation also requires a mechanism for trust calibration. This serves a dual purpose: ensuring that user feedback on classifications is reliable and providing the AI model with higher-quality additional training data. This trust-calibration process is integrated into the “4. On-Demand Editable Explanations” and “5. Counter-Examples” panels, as depicted in 1(b). Specifically, user feedback on explanations helps refine the AI model in different ways. If there is a direct, one-to-one relationship between an explanation and the model’s decision-making process—as seen in decision trees—the model can be updated immediately. In contrast, when using more complex models, such as deep learning, the feedback is transformed into additional data points that contribute to adjusting the model’s internal parameters.
4 Positioning Explanation-Driven Interventions as an EUD Tool for AI
The Rhino-Cyt intervention UI introduces a novel approach to EUD for AI, using editable explanations as an AI customization mechanism. This section explores Rhino-Cyt’s classification within the current EUD for AI landscape, comparing it with existing AI customization approaches while also examining its impact on human-AI symbiosis.
4.1 Main Features of the Intervention-Based Approach
Rhino-Cyt’s main features are summarized in 1 and described below to highlight the customization aspects of the interaction workflow.
Table 1: A summary of the main features of Rhino-Cyt’s intervention-based UI
Feature Rhino-Cyt’s Intervention UI
Customization Approach Explanation-driven intervention
User Control Freeform modifications of AI outputs and justifications
Impact on Model Behavior Direct, real-time adaptation
Technical Expertise Required None (domain expertise only)
Rhino-Cyt enables end users to customize its functionalities through explanation-driven intervention, providing insights into the decision-making process and identifying key aspects where user intervention is needed. Based on the information highlighted in the explanation, users can take actions such as marking the explanation as accurate or incorrect or adjusting the feature values used in the reasoning process. This approach puts physicians in control by allowing them to refine the system’s outputs and justifications, enhancing its performance, especially in cases involving outliers or exceptional situations. In this context, user actions can represent feedback for the system that can adapt its behavior over time. This feedback-driven process can lead physicians to directly intervene on the system’s behavior, fostering a real-time adaptation. By integrating these features, Rhino-Cyt can be considered a significant support tool for physicians, enabling them to perform their daily tasks more efficiently without requiring technical expertise to use the system.
4.2 Mapping Explanation-Driven Interventions to EUD for AI Taxonomies
To classify Rhino-Cyt within the existing EUD for AI landscape, we adopt a recent framework proposed by Esposito et al. [11], which categorizes EUD AI solutions based on the dimensions presented in 2.
Table 2: Classification of the intervention-based UI according to [11]
EUD Dimension Rhino-Cyt Implementation
Composition Paradigm Explanation-driven, rule-based intervention
Target Users Domain experts (rhinocytologists)
Technology AI-assisted medical diagnostics
Usage Single-user, with potential for collaborative interventions
Customization Level Tailoring and indirect model refinement
Approach Output AI model adaptation via explanation modifications
The composition paradigm of Rhino-Cyt provides explanation-driven intervention, meaning users can modify the AI model’s behavior through explanations of its decisions based on rules that the system follows to manage information. The system can be considered an AI-assisted medical diagnostic tool whose usage and functionalities are addressed to domain experts (i.e. rhinocytologists) who can employ this technology to improve the diagnostic process. The system is designed to potentially support collaborative usage in which multiple users can state their own points of view at the same time. Actually, the system supports only single-user usage, which allows tailoring, indirectly refining and adapting the model’s behavior to user needs through the features described in 4.1.
4.3 Comparison with Existing AI Customization Approaches
To illustrate the novelty of Rhino-Cyt, we compare it with three common AI customization paradigms: rule-based customization, no-code model building, and collaboration interfaces. The comparison is reported in 3
Table 3: Comparison between the three common customization paradigms and the Rhino-Cyt intervention-based UI.
Approach Customization Scope Technical Expertise Required Impact on AI Model
Rule-Based AI Customization Predefined rule sets Moderate Direct, deterministic changes
No-Code AI Model Building Component-based visual programming Low Configures AI before deployment
Human-AI Collaboration Interfaces Users validate/override AI outputs None No direct AI adaptation
Explanation-Driven Intervention Editable explanations influence AI reasoning None Indirect, adaptive refinements over time
3 shows how the Explanation-Driven Intervention enables users to manipulate the reasoning of the AI model based on the modification of the explanations. This approach becomes particularly effective in the medical context because no technical expertise is required: professionals can redirect and refine the systems’ behavior merely by relying on their knowledge and background in their field. As opposed to other approaches that either do not allow model adaptation or only modify it post-deployment, Rhino-Cyt facilitates a continuous and progressive adaptation over time through this interaction mechanism. Rhino-Cyt offers a more flexible adaptation process compared to rule-based AI customization, allowing experts to iteratively refine AI explanations rather than defining rigid rules, which enables adaptive learning. Unlike no-code AI development tools, Rhino-Cyt supports post-deployment model refinement, addressing the need for continuous AI adjustment in medical diagnostics. Additionally, while traditional human-AI collaboration interfaces rely on validation-based user feedback, Rhino-Cyt facilitates explanation-driven AI learning, ensuring the model evolves in alignment with expert reasoning.
4.4 Impact on Human-AI Collaboration and Trust Calibration
One of the biggest challenges in AI-assisted decision support is trust calibration, which is a well-documented issue in AI research, especially in high-stakes fields like medicine. More specifically, AI systems must be created, ensuring that users do not simply accept AI recommendations blindly—or, conversely, dismiss them outright. When users do not understand how a model achieves its conclusions, they might overtrust it, assuming it’s always correct, or undertrust it, ignoring useful insights out of skepticism. Calibrating trust is crucial to ensure that human expertise remains critical throughout the interaction while establishing the proper mechanisms to ensure their understanding of outputs and decisions [3, 19, 8]. The Rhino-Cyt intervention UI tackles this challenge by giving users meaningful ways to engage with the AI’s reasoning process rather than just its final outputs. First, it allows users to correct AI-generated explanations, ensuring that justifications align with their medical expertise. Instead of simply overriding a classification, they can refine the reasoning behind it, fostering a symbiotic relationship in which both the user and the AI system learn in the process [10]. Second, the system provides an “impact visualization dashboard”, giving users real-time feedback on how their interventions shape AI behavior over time. This transparency reinforces trust, showing that the AI isn’t a static black box but an adaptive tool that evolves based on expert input. Finally, Rhino-Cyt establishes a continuous feedback loop, where users’ refinements gradually steer the model that powers the system toward better, safer, and more reliable predictions [24]. Instead of treating AI as an inflexible system, this approach positions it as a collaborative partner, learning from domain expertise in a way that strengthens both accuracy and user confidence.
5 Conclusions and Future Work
This article introduced Rhino-Cyt as an EUD system for AI-assisted medical diagnostics, proposing explanation-driven intervention as a novel approach to AI customization. Traditional AI customization methods often limit the role of domain experts to passive reviewers. Rhino-Cyt, by contrast, enables experts to directly manipulate AI-generated outputs and their explanations, offering a way to refine AI behavior without requiring any programming expertise. Rhino-Cyt broadens the scope of AI customization and fosters trust calibration, ensuring that model outputs remain aligned with experts’ reasoning. Through its positioning within existing EUD taxonomies (2), this study highlighted Rhino-Cyt’s hybrid nature, sitting at the intersection of XAI and user-driven model customization. The three-step interaction workflow—classification overrides, explanation editing, and impact visualization—aims to allow domain experts to progressively adapt AI systems while maintaining meaningful control over decision-making processes. This work contributes to ongoing research in human-centered AI and AI-assisted decision support, demonstrating that intervention-based, explainability-driven EUD systems have the potential to bridge the gap between AI automation and expert oversight. By shifting AI-assisted diagnostics toward a model where human expertise plays an active role in shaping AI behavior, Rhino-Cyt paves the way for more transparent, trustworthy, and adaptable AI systems. The open questions regarding Rhino-Cyt and its implementation of explanation-driven interventions concern their usability, effectiveness, and broader applicability. The first step will be conducting a study with users (i.e., rhinocytologists) to assess its usability and determine its strengths and weaknesses to improve in the next development iteration. One key direction for future research is expanding its capabilities beyond single-user interactions. Medical diagnostics often rely on expert consensus, and extending Rhino-Cyt to support multi-user collaboration could enable a more robust AI adaptation process. Shared knowledge bases and federated learning could further enhance this by allowing explanation refinements to inform AI behavior across multiple institutions, ensuring that AI systems continuously learn from diverse expert insights. Another critical area of investigation is the impact of editable explanations on AI performance. Future studies should assess how user interventions influence model accuracy, interpretability, and cognitive load. A structured evaluation with domain experts could provide insights into whether explanation-driven interventions foster greater trust and understanding compared to traditional rule-based AI customization. Beyond the medical domain, the principles behind Rhino-Cyt could be applied to other fields, such as legal and financial AI, where professionals must interpret and refine AI-generated reasoning. As AI continues to integrate into high-stakes decision-making, ensuring that users can meaningfully shape its outputs will be essential for fostering reliable, human-centered AI systems. {credits}
5.0.1 Acknowledgements
This research is partially supported by the co-funding of the European Union - Next Generation EU: NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centers, companies and research D.D. MUR n. 341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research – FAIR” - CUP: H97G22000210007). This research is also partially supported by the Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” (Grant P2022JJPBY) - CUP: H53D23008150001. The research of Andrea Esposito is funded by a Ph.D. fellowship within the framework of the Italian “D.M. n. 352, April 9, 2022”- under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 – Ph.D. Project “Human-Centred Artificial Intelligence (HCAI) techniques for supporting end users interacting with AI systems,” co-supported by “Eusoft S.r.l.” (CUP H91I22000410007).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@
 Abstract
-[…]
-[name=Andrea Esposito, color=red]AE abstract The integration of Artificial Intelligence (AI) in modern society is heavily shifting the way that individuals carry out their tasks and activities.
-Employing AI-based systems raises challenges that designers and developers must address to ensure that humans remain in control of the interaction process, particularly in high-risk domains.
-This article presents a novel End-User Development (EUD) approach for black-box AI models through a redesigned user interface in the Rhino-Cyt platform, a medical AI-based decision-support system for medical professionals (more precisely, rhinocytologists) to carry out cell classification.
-The proposed interface empowers users to intervene in AI decision-making process by editing explanations and reconfiguring the model, influencing its future predictions.
-This work contributes to Human-Centered Artificial Intelligence (HCAI) and EUD by discussing how explanation-driven interventions allow a blend of explainability, user intervention, and model reconfiguration, fostering a symbiosis between humans and user-tailored AI systems.
+The integration of Artificial Intelligence (AI) in modern society is transforming how individuals perform tasks.
+In high-risk domains, ensuring human control over AI systems remains a key design challenge.
+This article presents a novel End-User Development (EUD) approach for black-box AI models, enabling users to edit explanations and influence future predictions through targeted interventions.
+By combining explainability, user control, and model adaptability, the proposed method advances Human-Centered AI (HCAI), promoting a symbiotic relationship between humans and adaptive, user-tailored AI systems.
 Keywords:
 Customization Black-box AI Model Reconfiguration Explainable AI (XAI) Human-AI symbiosis
-abstract The integration of Artificial Intelligence (AI) in modern society is heavily shifting the way that individuals carry out their tasks and activities.
-Employing AI-based systems raises challenges that designers and developers must address to ensure that humans remain in control of the interaction process, particularly in high-risk domains.
-This article presents a novel End-User Development (EUD) approach for black-box AI models through a redesigned user interface in the Rhino-Cyt platform, a medical AI-based decision-support system for medical professionals (more precisely, rhinocytologists) to carry out cell classification.
-The proposed interface empowers users to intervene in AI decision-making process by editing explanations and reconfiguring the model, influencing its future predictions.
-This work contributes to Human-Centered Artificial Intelligence (HCAI) and EUD by discussing how explanation-driven interventions allow a blend of explainability, user intervention, and model reconfiguration, fostering a symbiosis between humans and user-tailored AI systems.
 EUD
@@
 1 Introduction
-AI has become an integral component of decision-support systems in numerous domains, including healthcare, finance, and law [15].
-While AI models could increase efficiency and decision-making capabilities, their reliance on complex, often opaque, algorithms presents a significant barrier to adoption [15].
-In high-stakes applications such as medical diagnostics, end-users—typically domain experts rather than AI specialists—require mechanisms to refine and adjust AI behavior to better align with their expertise and contextual knowledge [8].
-Traditionally, AI systems have followed a one-size-fits-all approach, where models are trained on large datasets but offer limited opportunities for users to modify their behavior post-deployment [7].
+Artificial Intelligence (AI) has become an integral component of decision-support systems in numerous domains, including healthcare, finance, and law [16].
+While AI models could enhance decision-making, their reliance on complex, often opaque, algorithms presents a significant barrier to adoption [16], especially in high-stakes applications such as medical diagnostics [6].
+End-users—typically domain experts rather than AI specialists—require mechanisms to refine and adjust AI behavior to better align with their expertise and contextual knowledge.
+Most AI systems follow a one-size-fits-all approach, with limited support for post-deployment customization [5].
 This lack of adaptability can lead to misaligned recommendations, loss of trust, and decreased usability.
-EUD for AI seeks to address this challenge by enabling non-technical users to customize AI behavior according to their specific needs and preferences [12].
-However, existing EUD approaches primarily focus on low-code or no-code AI development environments, offering component-based or rule-based interactions [18].
-Few solutions have explored explanation-driven interventions, where users influence AI behavior by modifying its justifications rather than its internal mechanics.
-The integration of EUD for AI has made substantial progress in fields such as the Internet of Things (IoT), education, and business analytics [11, 18].
-However, AI-based decision-support systems, especially those powered by black-box models, remain challenging for users to directly intervene on.
+End-User Development (EUD) for AI seeks to address this issue by enabling non-technical users to customize AI behavior according to their specific needs [12].
+While the integration of EUD for AI has made substantial progress in fields such as the Internet of Things (IoT), education, and business analytics [11, 20], AI-based decision-support systems—especially those powered by black-box models—remain challenging for users to modify.
 Current approaches to AI customization for end-users are (i) Rule-Based Customization[11], (ii) Low-Code / No-Code AI[20], and (iii) Human-AI Collaboration Interfaces[8, 22].
 The first one involves the definition from the user of if-this-then-that conditions to influence AI outputs.
-Although effective in structured environments, this approach lacks flexibility in complex decision-making scenarios.
 On the other hand, the platforms that adopt Low-Code / No-Code AI allow users to build and deploy AI models without programming (e.g., AutoML tools), but they do not enable real-time intervention on model behavior post-deployment.
 Lastly, Human-AI Collaboration Interfaces enable users to validate and/or override the system’s predictions without affecting its future behavior.
-These approaches do not fully address the need for an interactive and iterative refinement of AI behavior based on human expertise, which can be highly useful in critical fields.
-For example, medical professionals might encounter AI misclassifications that can be reviewed and corrected through targeted feedback.
-Although the latter is usually employed to merely accept or reject the output without repercussions on the model, using professionals’ knowledge and expertise to refine AI reasoning can be a valuable resource to improve the system’s performance, increase accuracy for future predictions and build a stronger symbiotic relationship between the two parties [14, 10].
-Creating AI-based systems that embody these characteristics can foster collaboration, which is central in high-stakes domains.
-Thus, end-users must be provided with clear, appropriate, and effective interaction mechanisms that enable bidirectional communication, establishing a symbiotic relationship between humans and AI.
+These approaches do not fully address the need for an interactive and iterative refinement of AI behavior based on human expertise.
+Using professionals’ knowledge and expertise to refine AI reasoning can be a valuable resource to improve the system’s performance while building a stronger symbiotic relationship between the two parties [15, 8].
+Creating AI-based systems that embody these characteristics can foster collaboration, establishing a symbiotic relationship between humans and AI.
 Symbiotic Artificial Intelligence (SAI) is a specialization of Human-Centered AI [26] and aims at supporting humans instead of replacing them.
 This implies creating solutions that reflect humans’ needs and preferences by integrating intervention paradigms, transparency, and fairness by design focusing on augmentation rather than automation [17, 8].
-This research presents a novel intervention-based User Interface (UI) within the Rhino-Cyt platform, designed to empower rhinocytologists to modify AI-generated classifications and explanations.
-Rhino-Cyt is an AI-assisted EUD environment for the classification of nasal cytology samples that supports its end-user developers, i.e., medical professionals, in diagnosing inflammatory and allergic conditions [2].
-Rhino-Cyt aims at embodying the EUDability construct [5] that introduces an innovative explanation-driven EUD approach, allowing end-users to adjust AI classifications, edit AI-generated explanations, and indirectly refine and tailor the AI model through the mechanism of interventions [22].
+This research proposes a novel intervention-based User Interface (UI) within the Rhino-Cyt platform, designed to empower rhinocytologists to modify AI-generated classifications and explanations.
+This proposal aims at embodying the EUDability construct [4] by introducing an innovative explanation-driven EUD approach, allowing end-users to adjust AI classifications, edit AI-generated explanations, and indirectly refine and tailor the AI model through the mechanism of interventions [24].
 This approach goes beyond rule-based or component-based customization, offering a human-in-the-loop model refinement paradigm.
@@
 The rest of the article is structured as follows.
-2 discusses prior research in EUD for AI, explainability, and human-AI collaboration, positioning Rhino-Cyt within this landscape.
-3 presents the design of the intervention-based UI, detailing its interaction flow and impact on AI adaptation.
-4 positions Rhino-Cyt within existing EUD for AI taxonomies and compares it with other customization paradigms.
-5 concludes this article by summarizing its key contributions and outlining the next steps.
+Section 2 discusses prior research in EUD for AI, explainability, and human-AI collaboration, also presenting Rhino-Cyt.
+Section 3 presents the design of the intervention-based UI, detailing its interaction flow and impact on AI adaptation.
+Section 4 positions Rhino-Cyt within existing EUD for AI taxonomies and compares it with other customization paradigms.
+Section 5 concludes this article by summarizing its key contributions and outlining the next steps.
 2 Background and Related Work
 This section reviews prior research on EUD for AI, explainability as a mechanism for EUD, and human-AI collaboration in decision-support systems.
-In the end, we highlight the explanation-driven intervention paradigm for AI-assisted medical diagnostics as the main contribution of this work.
 2.1 End-User Development for AI
 EUD encompasses a range of methods, techniques, tools, and socio-technical environments that empower non-professionals to engage in activities usually reserved for professionals in ICT-related areas, including the ability to create, modify, extend, and test digital artifacts without requiring specialized knowledge in conventional software engineering practices [3].
-In the context of AI, EUD plays a crucial role in enabling non-technical users to adjust AI systems without requiring programming expertise.
 A systematic literature review by Esposito et al.
@@
     Workflow and Data Diagrams: Users define AI processes using flow-based representations.
-Most EUD for AI systems fall within component-based and rule-based paradigms, where users interact with structured representations of AI logic [11].
 While these approaches are effective for tasks such as building AI models from scratch or configuring predefined automation rules, they offer limited support for modifying existing black-box AI models.
-The Rhino-Cyt intervention interface extends EUD for AI by introducing a new paradigm: explanation-driven interventions.
-Instead of requiring users to manipulate AI model components or logic directly, this approach allows them to edit AI-generated explanations, indirectly refining the model’s behavior over time.
-This method aligns with the goal of making AI more adaptable to domain-specific knowledge while minimizing technical barriers.
-2.2 Explainability as a Mechanism for End-User Development
+Explanation-driven interventions extend EUD for AI by introducing a new paradigm.
+Instead of requiring users to manipulate AI model components or logic directly, this proposal allows them to edit AI-generated explanations, indirectly refining the model’s behavior over time.
 […]
-XAI seeks to make AI model decisions more interpretable and transparent, particularly for domain experts who rely on AI assistance in high-stakes decision-making scenarios [11, 21].
-Traditionally, eXplainable Artificial Intelligence (XAI) was mostly used for post-hoc justification, trust calibration, and bias detection.
-Through XAI, users can understand AI decisions, deciding whether to rely on its predictions or potentially recognize (and mitigate) biases by assessing its reasoning [15].
-However, explainability has rarely been explored as an active mechanism for EUD.
-In most AI-assisted decision systems, explanations are static: users can view and interpret them, but they cannot modify them to influence future AI behavior.
-Rhino-Cyt introduces a novel editable explanation mechanism where users can: modify explanations associated with AI classifications, provide domain-specific refinements to ensure AI-generated explanations align with expert knowledge, and influence future AI behavior indirectly, fostering an interactive trust-building process.
+In most AI-assisted decision systems, explanations are static, not allowing users to modify them to influence future AI behavior.
+Our proposal introduces a novel editable explanation mechanism where users can modify explanations associated with AI classifications, provide domain-specific refinements to ensure AI-generated explanations align with expert knowledge, and influence future AI behavior indirectly.
 2.2 Human-AI Collaboration in Decision Support Systems
@@
 This translates into a validation-based collaboration, where experts review AI predictions but have no direct manipulation mechanism for modifying the AI’s reasoning process [8].
-Rhino-Cyt aims at filling this gap following the model of human–AI interaction proposed by Desolda et al.
+Our proposal aims at filling this gap following the model of human–AI interaction proposed by Desolda et al.
 [7], moving beyond this by enabling direct, explanation-driven interventions, ensuring that AI’s decision-making processes evolve alongside domain experts, fostering a symbiosis among humans and AI [8].
-Establishing proper collaboration mechanisms between humans and AI is crucial to guarantee that professionals are aware of the processes that lead to outputs; in fact, the outcome of the interaction must be a fair, non-discriminatory, and unbiased decision that minimizes potentially negative repercussions and holds accountable both parties.
-At the same time, humans remain the ones responsible for perpetrating the decision in their application domain [25, 6].
+Establishing proper collaboration mechanisms between humans and AI is crucial to guarantee that professionals are aware of the processes that lead to outputs.
+2.3 Rhino-Cyt: an AI-based system for Rhinocitology
+Our proposal leverages a case study in the context of rhinocytology (a subfield of medical cytology) [9, 10].
+Currently, the diagnostic process in rhinocytology is mainly based on direct observation under the microscope, which requires a prolonged effort by rhinocytologists [14, 7].
+Rhino-Cyt is an AI-assisted environment for the classification of nasal cytology samples that supports its end-user developers, i.e., medical professionals, in diagnosing inflammatory and allergic conditions [10].
+It employs AI models (namely, a CNN) to automate the cytological examination by segmenting histological samples of the nasal mucosa, identifying and classifying individual cells based on nine cytotypes [10, 14]: (i) ciliated, (ii) muciparous, (iii) basal cells, (iv) striated cells, (v) neutrophils, (vi) eosinophils, (vii) mast cells, (viii) lymphocytes, (ix) metaplastic cells.
 3 Design of the Intervention-Based User Interface
-(a) The intervention-based UI
-(b) Key areas of the UI
+Rhino-Cyt introduces an innovative intervention-based UI that allows rhinocytologists to refine AI-assisted classifications, presented in Fig.
+1.
+Users can intervene modifying decisions and explanations, unlike conventional AI customization methods, which rely on rule-based or low-code paradigms [24].
 Figure 1: The user interface for allowing black-box AI model tailoring implemented in Rhino-Cyt
-The Rhino-Cyt platform introduces an innovative intervention-based UI that allows rhinocytologists to refine AI-assisted nasal cytology classifications, presented in 1.
-Unlike conventional AI customization methods, which rely on rule-based or low-code paradigms, Rhino-Cyt allows users to intervene in AI decisions and modify AI-generated explanations, indirectly influencing the model’s future behavior [22].
 This section details the design principles, UI components, and interaction workflow, demonstrating how the system aligns with EUD for AI by offering an explanation-driven intervention mechanism.
@@
   1. 1.
-    Intervention-Based Interaction.
-The system enables users to adjust AI-generated classifications and explanations in a structured manner.
-Any modifications made by users are systematically logged and tracked, allowing for a comprehensive assessment of their impact on the AI model’s behavior over time.
+    Intervention-Based Interaction. The system allows logged users to adjust AI classifications and explanations tracking any modifications.
   2. 2.
     Explainability-Driven Customization.
-The system supports explainability-driven customization by allowing users to refine AI-generated justifications without requiring them to modify raw model parameters or write explicit rules.
+The system supports it by allowing users to refine AI-generated justifications without requiring them to modify raw model parameters or write explicit rules.
 Instead, the interface leverages editable explanations as a means of customization, enabling domain experts to adjust and improve AI-generated reasoning based on their medical expertise.
@@
 3.2 User Interface Components
-The elements of Rhino-Cyt’s UI that allow users to intervene on AI decisions consist of three key components:
+The elements of Rhino-Cyt’s UI (Fig. 1) that allow users to intervene on AI decisions consist of the following key components
   1. 1.
-    The Classification Details and Interventions Panel presents AI-generated classifications for nasal cytology samples, allowing users to review system decisions.
-If necessary, users can override AI’s decisions by selecting an alternative category, ensuring that domain expert judgment remains central to the decision-making process.
-Any corrections made by users are logged and incorporated into the system, contributing to the continuous adaptation and improvement of the AI’s future performance.
+    The Classification Details and Interventions Panel presents AI-generated classifications, allowing users to review system decisions.
+If necessary, users can override AI’s decisions by selecting an alternative category which is incorporated into the system, contributing to the continuous adaptation and improvement of the AI’s future performance.
   2. 2.
-    The Editable Explanation Area displays the AI-generated rationale behind its classifications, providing transparency into the decision-making process.
-Users can modify these explanations to guarantee they align more closely with medical reasoning and domain expertise.
+    The Editable Explanation Area displays the AI-generated justifications of its classifications, providing transparency into the decision-making process.
+Users can modify these justifications to guarantee they align more closely with their expertise.
 Any edits made by users directly influence the AI’s reasoning model, refining its approach and shaping future justifications.
-  3. 3.
-    The Impact Visualization Dashboard provides a summary of user interventions, detailing key aspects such as the number of classification overrides, the frequency and nature of explanation edits, and the ways in which AI predictions evolve over time based on these interventions.
-This feature plays a crucial role in trust calibration by offering users insights into how the AI adapts and improves, helping users to understand how AI behavior is adapting.
-By combining those, Rhino-Cyt enables domain experts to refine the AI decision-making process by directly manipulating its output, without requiring technical expertise.
 3.3 Interaction Workflow: How Users Intervene in AI Decisions
-The Rhino-Cyt intervention workflow is structured as a three-step process, ensuring smooth user interaction with the AI system.
+The Rhino-Cyt intervention workflow is structured as a three-step process.
   1. Step 1:
-    Reviewing AI-Generated Classification and Explanation.
-Upon analyzing a nasal cytology sample, the AI presents a predicted classification along with a textual explanation justifying the classification.
+    Reviewing AI-Generated Classification and Explanation. Upon analyzing a nasal cytology sample, the AI presents a predicted classification and a textual explanation justifying the classification.
   2. Step 2:
@@
 The users have two options for intervention: they can either override the AI classification by selecting an alternative label or modify the AI-generated explanation to reflect expert reasoning more accurately, or both.
-Any changes are recorded, with an optional comment field for contextualizing edits.
   3. Step 3:
-    Model Adaptation and Visualization of Impact. The system logs interventions and updates the AI’s explanation model allowing users to track how their interventions can shape future AI predictions.
+    Model Adaptation and Visualization of Impact. The system logs interventions and updates the AI’s explanation model.
 This workflow enables AI adaptation to be progressive, ensuring the model evolves alongside experts’ knowledge, thus creating a human-in-the-loop customization mechanism for AI-supported decision-making.
 3.4 Underlying AI Model Adaptation
-Rhino-Cyt employs a hybrid adaptation mechanism.
-It balances manual user intervention with an automated model refinement process based on users’ feedback, thus providing an example of both an adaptable and adaptive system [12].
-The system incremental adaptation is powered by three key requirements: (i) interaction logging, (ii) a quick AI training loop, and (iii) a trust-calibration mechanism.
+Our proposal balances manual user intervention with an automated model refinement process based on users’ feedback, thus providing an example of both an adaptable and adaptive system [12].
 The platform continuously logs user interactions, capturing every action performed by users.
 This includes explicit feedback on the AI model’s predictions, where users can either accept the suggested classification or reject it and make modifications.
-This process is facilitated through the “3.
-Intervention” panel, illustrated in 1(b).
 The collected feedback is then leveraged to initiate the retraining of the AI model.
-To enhance the adaptivity and adaptability of the AI system, several techniques can be employed, ranging from online learning [1, 26] to reinforcement learning from human feedback [17], ensuring timely updates to the model.
-However, effective adaptation also requires a mechanism for trust calibration.
-This serves a dual purpose: ensuring that user feedback on classifications is reliable and providing the AI model with higher-quality additional training data.
-This trust-calibration process is integrated into the “4.
-On-Demand Editable Explanations” and “5.
-Counter-Examples” panels, as depicted in 1(b).
+To enhance the adaptivity and adaptability of the AI system, several techniques can be employed, ranging from online learning [1, 27] to reinforcement learning from human feedback [18], leveraging detailed explanation obtained either from a gray-box model (as the one presented by Desolda et al.
+[7]) or through large-language models.
 Specifically, user feedback on explanations helps refine the AI model in different ways.
@@
 This section explores Rhino-Cyt’s classification within the current EUD for AI landscape, comparing it with existing AI customization approaches while also examining its impact on human-AI symbiosis.
-4.1 Main Features of the Intervention-Based Approach
-Rhino-Cyt’s main features are summarized in 1 and described below to highlight the customization aspects of the interaction workflow.
+Rhino-Cyt’s main features are summarized in Table 1 and described below to highlight the customization aspects of the interaction workflow.
 Table 1: A summary of the main features of Rhino-Cyt’s intervention-based UI
@@
 Technical Expertise Required None (domain expertise only)
-Rhino-Cyt enables end users to customize its functionalities through explanation-driven intervention, providing insights into the decision-making process and identifying key aspects where user intervention is needed.
+It enables end-users to customize its functionalities through explanation-driven intervention.
 Based on the information highlighted in the explanation, users can take actions such as marking the explanation as accurate or incorrect or adjusting the feature values used in the reasoning process.
-This approach puts physicians in control by allowing them to refine the system’s outputs and justifications, enhancing its performance, especially in cases involving outliers or exceptional situations.
+This approach puts physicians in control by allowing them to refine the system’s outputs and justifications, enhancing its performance, especially in cases involving outliers.
 In this context, user actions can represent feedback for the system that can adapt its behavior over time.
-This feedback-driven process can lead physicians to directly intervene on the system’s behavior, fostering a real-time adaptation.
-By integrating these features, Rhino-Cyt can be considered a significant support tool for physicians, enabling them to perform their daily tasks more efficiently without requiring technical expertise to use the system.
-4.2 Mapping Explanation-Driven Interventions to EUD for AI Taxonomies
-To classify Rhino-Cyt within the existing EUD for AI landscape, we adopt a recent framework proposed by Esposito et al. [11], which categorizes EUD AI solutions based on the dimensions presented in 2.
+To classify Rhino-Cyt within the existing EUD for AI landscape, we adopt a recent framework proposed by Esposito et al.
+[11], which categorizes EUD AI solutions based on the dimensions presented in Table 2.
 Table 2: Classification of the intervention-based UI according to [11]
@@
 Approach Output AI model adaptation via explanation modifications
-The composition paradigm of Rhino-Cyt provides explanation-driven intervention, meaning users can modify the AI model’s behavior through explanations of its decisions based on rules that the system follows to manage information.
-The system can be considered an AI-assisted medical diagnostic tool whose usage and functionalities are addressed to domain experts (i.e.
-rhinocytologists) who can employ this technology to improve the diagnostic process.
-The system is designed to potentially support collaborative usage in which multiple users can state their own points of view at the same time.
-Actually, the system supports only single-user usage, which allows tailoring, indirectly refining and adapting the model’s behavior to user needs through the features described in 4.1.
-4.3 Comparison with Existing AI Customization Approaches
 To illustrate the novelty of Rhino-Cyt, we compare it with three common AI customization paradigms: rule-based customization, no-code model building, and collaboration interfaces.
-The comparison is reported in 3
+The comparison is reported in Table 3.
 Table 3: Comparison between the three common customization paradigms and the Rhino-Cyt intervention-based UI.
@@
 Explanation-Driven Intervention Editable explanations influence AI reasoning None Indirect, adaptive refinements over time
-3 shows how the Explanation-Driven Intervention enables users to manipulate the reasoning of the AI model based on the modification of the explanations.
+Explanation-Driven Intervention enables users to manipulate the reasoning of the AI model by modifying the explanations.
 This approach becomes particularly effective in the medical context because no technical expertise is required: professionals can redirect and refine the systems’ behavior merely by relying on their knowledge and background in their field.
-As opposed to other approaches that either do not allow model adaptation or only modify it post-deployment, Rhino-Cyt facilitates a continuous and progressive adaptation over time through this interaction mechanism.
-Rhino-Cyt offers a more flexible adaptation process compared to rule-based AI customization, allowing experts to iteratively refine AI explanations rather than defining rigid rules, which enables adaptive learning.
-Unlike no-code AI development tools, Rhino-Cyt supports post-deployment model refinement, addressing the need for continuous AI adjustment in medical diagnostics.
-Additionally, while traditional human-AI collaboration interfaces rely on validation-based user feedback, Rhino-Cyt facilitates explanation-driven AI learning, ensuring the model evolves in alignment with expert reasoning.
-4.4 Impact on Human-AI Collaboration and Trust Calibration
-One of the biggest challenges in AI-assisted decision support is trust calibration, which is a well-documented issue in AI research, especially in high-stakes fields like medicine.
-More specifically, AI systems must be created, ensuring that users do not simply accept AI recommendations blindly—or, conversely, dismiss them outright.
-When users do not understand how a model achieves its conclusions, they might overtrust it, assuming it’s always correct, or undertrust it, ignoring useful insights out of skepticism.
-Calibrating trust is crucial to ensure that human expertise remains critical throughout the interaction while establishing the proper mechanisms to ensure their understanding of outputs and decisions [3, 19, 8].
-The Rhino-Cyt intervention UI tackles this challenge by giving users meaningful ways to engage with the AI’s reasoning process rather than just its final outputs.
-First, it allows users to correct AI-generated explanations, ensuring that justifications align with their medical expertise.
-Instead of simply overriding a classification, they can refine the reasoning behind it, fostering a symbiotic relationship in which both the user and the AI system learn in the process [10].
-Second, the system provides an “impact visualization dashboard”, giving users real-time feedback on how their interventions shape AI behavior over time.
-This transparency reinforces trust, showing that the AI isn’t a static black box but an adaptive tool that evolves based on expert input.
-Finally, Rhino-Cyt establishes a continuous feedback loop, where users’ refinements gradually steer the model that powers the system toward better, safer, and more reliable predictions [24].
-Instead of treating AI as an inflexible system, this approach positions it as a collaborative partner, learning from domain expertise in a way that strengthens both accuracy and user confidence.
+This approach enables adaptive learning ensuring the model evolves in alignment with expert reasoning.
+Trust calibration is a major challenge in AI-assisted decision support, particularly in high-stakes domains like medicine [19].
+More specifically, AI systems must be created ensuring that users do not simply accept AI recommendations blindly—or, conversely, dismiss them outright [2, 21, 6].
+Our proposal tackles this challenge by allowing users to engage with the AI’s reasoning process rather than just its final outputs.
+Instead of simply overriding a classification, experts can refine the reasoning behind it, fostering a symbiotic relationship in which both the user and the AI system learn in the process [8].
+Through explanation-driven interventions, users’ refinements gradually steer the AI model toward better, safer, and more reliable predictions.
+This approach positions AI as a collaborative partner, learning from domain expertise in a way that strengthens both accuracy and user confidence.
 5 Conclusions and Future Work
-This article introduced Rhino-Cyt as an EUD system for AI-assisted medical diagnostics, proposing explanation-driven intervention as a novel approach to AI customization.
-Traditional AI customization methods often limit the role of domain experts to passive reviewers.
-Rhino-Cyt, by contrast, enables experts to directly manipulate AI-generated outputs and their explanations, offering a way to refine AI behavior without requiring any programming expertise.
-Rhino-Cyt broadens the scope of AI customization and fosters trust calibration, ensuring that model outputs remain aligned with experts’ reasoning.
-Through its positioning within existing EUD taxonomies (2), this study highlighted Rhino-Cyt’s hybrid nature, sitting at the intersection of XAI and user-driven model customization.
-The three-step interaction workflow—classification overrides, explanation editing, and impact visualization—aims to allow domain experts to progressively adapt AI systems while maintaining meaningful control over decision-making processes.
-This work contributes to ongoing research in human-centered AI and AI-assisted decision support, demonstrating that intervention-based, explainability-driven EUD systems have the potential to bridge the gap between AI automation and expert oversight.
-By shifting AI-assisted diagnostics toward a model where human expertise plays an active role in shaping AI behavior, Rhino-Cyt paves the way for more transparent, trustworthy, and adaptable AI systems.
-The open questions regarding Rhino-Cyt and its implementation of explanation-driven interventions concern their usability, effectiveness, and broader applicability.
-The first step will be conducting a study with users (i.e., rhinocytologists) to assess its usability and determine its strengths and weaknesses to improve in the next development iteration.
-One key direction for future research is expanding its capabilities beyond single-user interactions.
-Medical diagnostics often rely on expert consensus, and extending Rhino-Cyt to support multi-user collaboration could enable a more robust AI adaptation process.
-Shared knowledge bases and federated learning could further enhance this by allowing explanation refinements to inform AI behavior across multiple institutions, ensuring that AI systems continuously learn from diverse expert insights.
-Another critical area of investigation is the impact of editable explanations on AI performance.
-Future studies should assess how user interventions influence model accuracy, interpretability, and cognitive load.
-A structured evaluation with domain experts could provide insights into whether explanation-driven interventions foster greater trust and understanding compared to traditional rule-based AI customization.
-Beyond the medical domain, the principles behind Rhino-Cyt could be applied to other fields, such as legal and financial AI, where professionals must interpret and refine AI-generated reasoning.
-As AI continues to integrate into high-stakes decision-making, ensuring that users can meaningfully shape its outputs will be essential for fostering reliable, human-centered AI systems.
+This article proposed explanation-driven interventions as an EUD tool for black-box AI systems, illustrated through their adoption in Rhino-Cyt, a system aiding rhinocytologists in cell counting tasks.
+Unlike traditional AI customization methods that relegate domain experts to passive reviewers, explanation-driven interventions empowers them to directly manipulate AI-generated outputs and explanations, without any programming expertise.
+This work contributes to ongoing research in human-centered AI and AI-assisted decision supports EUD systems have the potential to bridge the gap between AI automation and expert oversight.
+Future work includes evaluating the usability and effectiveness of Rhino-Cyt through user studies, identifying improvement areas, and extending support to multi-user collaboration.
+Better methods for tracking intervention impact and mitigating user errors are also needed.
+Another key direction is assessing how editable explanations affect model performance, accuracy, and user cognitive load.
+Structured evaluations could clarify whether this approach enhances trust and understanding more than traditional rule-based customization.
+Finally, explanation-driven interventions hold promise beyond medicine, with potential in domains like law and finance where professionals must interpret and refine AI reasoning.
 {credits}
 5.0.1 Acknowledgements
-This research is partially supported by the co-funding of the European Union - Next Generation EU: NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centers, companies and research D.D.
+This research is partially supported by: (i) the co-funding of the European Union - Next Generation EU: NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centers, companies and research D.D.
 MUR n.
-341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research – FAIR” - CUP: H97G22000210007).
-This research is also partially supported by the Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” (Grant P2022JJPBY) - CUP: H53D23008150001.
+341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research – FAIR” - CUP: H97G22000210007); (ii) the Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” (Grant P2022JJPBY) - CUP: H53D23008150001.
 The research of Andrea Esposito is funded by a Ph.D.
@@
 Project “Human-Centred Artificial Intelligence (HCAI) techniques for supporting end users interacting with AI systems,” co-supported by “Eusoft S.r.l.” (CUP H91I22000410007).
+The research of Francesco Greco is funded by a Ph.D.
+fellowship within the framework of the Italian “D.M.
+n.
+352, April 9, 2022”- under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 – Ph.D.
+Project “Investigating XAI techniques to help user defend from phishing attacks”, co-supported by “Auriga S.p.A.” (CUP H91I22000410007).
 […]